OpenStudy (summersnow8):

A U-tube with a cross-sectional area of 1.00 cm2 is open to the atmosphere at both ends. Water is poured into the tube until the water rises part-way along the straight sides, and then 5.00 cm3 of oil is poured into one end. As a result, the top surface of the oil ends up 0.550 cm higher than the water surface on the other side of the U. What is the density of the oil? Answer: _______ kg/m3

OpenStudy (shamim):

the density of oil is less than the density of water

OpenStudy (summersnow8):

5.39d

OpenStudy (shamim):

pressure of both sides r equal

OpenStudy (shamim):

i know thr density of water is 1gm/cm^3

OpenStudy (summersnow8):

well it is 1000kg/m^3

OpenStudy (shamim):

1000kg/m^3=1000*1000gm/10^6cm^3

OpenStudy (shamim):

=1gm/cm^3

OpenStudy (summersnow8):

okay, but we are looking for kg/m3

OpenStudy (summersnow8):

is this right? 5-0.550=4.45 5cm^3 of oil=4.45cm^3 of water density of oil=4.45/5=0.89 of water 0.89*1000=890km/m^3

OpenStudy (summersnow8):

@shamim

OpenStudy (shamim):

ya i think ur result is ok

OpenStudy (summersnow8):

so its 890kg/m^3

OpenStudy (shamim):

in my equation h= height of water

OpenStudy (summersnow8):

okay, so you believe that mine is correct? did you get the same as me?

OpenStudy (shamim):

h*1*9.8=(h-4.45)*1*9.8+5*d*9.8 h=height of water level g=9.8 m/sec^2 1gm/cm^3 is the density of watet

OpenStudy (shamim):

ya my result will b same as u

OpenStudy (shamim):

but my result is in cgs system we hv to convert it into mks system

OpenStudy (shamim):

just multiply by 1000

OpenStudy (shamim):

did u get my solution

OpenStudy (summersnow8):

hmm i dont understand how you solved it

OpenStudy (shamim):

i think mine is better for understanding

OpenStudy (summersnow8):

well can you show me the equation then show me the numbers plugged in, im really confused

OpenStudy (shamim):

ok

OpenStudy (shamim):

c my equation

OpenStudy (shamim):

u know the pressure of both sides r equal

OpenStudy (shamim):

one side is full of water

OpenStudy (shamim):

let the height of that water is h

OpenStudy (shamim):

other side is full of water n oil

OpenStudy (shamim):

water height is less than the water height of other side by 4.45cm

OpenStudy (shamim):

height of oil is 5 cm

OpenStudy (summersnow8):

im way confused :/

OpenStudy (shamim):

u know the equation of pressure is p=hdg h=height of liquid d=density of lequid g=9.8 m/sec^2

OpenStudy (shamim):

i hv to draw it for better understanding

OpenStudy (shamim):

i m now frm my cell phone so its not possible to dtaw

OpenStudy (shamim):

draw

OpenStudy (summersnow8):

but i am confused on what the value of h is and the density of the liquid....

OpenStudy (shamim):

i c

OpenStudy (shamim):

h will b cancell out

OpenStudy (shamim):

n we will find out the value of liquid d

OpenStudy (summersnow8):

okay, you have to write me the equation that gives me 890, because im not understanding how you got yours

OpenStudy (shamim):

u will take all h terms in one side

OpenStudy (shamim):

h*1*9.8=(h-4.45)*1*9.8+5*d*9.8 h=(h-4.45)+5*d i cancelled 9.8 frm each term

OpenStudy (shamim):

5*d=h-h+4.45

OpenStudy (shamim):

5*d=4.45

OpenStudy (shamim):

did u get now

OpenStudy (summersnow8):

no, i dont

OpenStudy (shamim):

where is ur problem

OpenStudy (shamim):

d=4.45/5 gm/ cm^3 d=(4.45/5)*1000 kg/m^3

OpenStudy (summersnow8):

okay

OpenStudy (shamim):

did u really get my solution